OpenStudy (anonymous):

Write the mixed number as a percent 2 3/50

OpenStudy (unklerhaukus):

\[206~\%\]

OpenStudy (anonymous):

how did u get the answer

OpenStudy (unklerhaukus):

well 1= 100% so 2=200% 3/50=6/100=6% therefore 2 and 3 fiftieths is 206%

OpenStudy (anonymous):

Thank you.